a sample of oxygen gas,which is saturated with water vapor,is kept in a 10 L vessel at 30°C and has a pressure of 758 Torr. If the pressure of the water vapour at this temperature is 31.8 what would be the pressure of the dry oxygen?


1
Expert's answer
2020-12-01T09:37:22-0500

p(total)=p1+p2.

p2=p(total)-p1 = 758-31.8 = 726,2 torr.